# Tropical curves of unibranch points and hypertangency

### v2

@@ 1.1. Results
In this paper we study unibranch points of curves, i.e. points which have a single preimage in the normalization of the curve. Our first goal is to study under what conditions two projective plane curves of degree at least $2$ intersect in exactly one point which is unibranch for both of them. This problem is also relevant for the study of the Lang exceptional set appearing in some conjectures related to arithmetic and hyperbolic geometry, as we shall explain later in this introduction.
@@
Let $B,C\subset\mathbb{P}^{2}$ be integral curves of degree at least $2$ such that $B\cap C=\{q\}$, where $q$ is an $(m^{\prime},n^{\prime})$-point for $B$ and an $(m,n)$-point for $C$. Then there is an isomorphism of tropical curves between $\Gamma^{(m,n)}$ and $\Gamma^{(m^{\prime},n^{\prime})}$; equivalently, $m^{\prime}/n^{\prime}=m/n$.
@@ 1.2. Context
Our work in this paper is motivated by the algebro-geometric counterparts of some outstanding conjectures from diophantine and hyperbolic geometry. To introduce them, let $B\subset\mathbb{P}^{2}$ be a complex projective curve having at most nodal singularities. The so-called algebraic exceptional set of $B$, introduced by Lang, is denoted as follows
@@
As an example of application of this type of results we mention that, in our previous paper [CT25], we used as a crucial ingredient a special case of Theorem 1.1.1 (Theorem 2.2.1 in loc.cit. which assumes $B$ is a smooth at $q$) to show that $\mathcal{E}(B)$ is empty for a general $B$ with three irreducible components of total degree at least 5.
@@ 1.3. Notation
Given a point $p\in C$, we write ${\operatorname{mult}}_{p}(C)$ for the multiplicity of $C$ at $p$; if $m={\operatorname{mult}}_{p}(C)$ we say that $p$ is an $m$-fold point . We denote by $\nu_{C}:C\to C$ the normalization. If $|\nu_{C}^{-1}(p)|=1$ we say that $p$ is unibranch.
@@ Acknowledgements.
We are grateful to Pietro Corvaja, Kristin DeVleming, Luca Ferrigno, Nicola Ottolini and Wei Chen for useful comments. Special thanks to Laura Capuano for showing us Lemma 5.1.2. LC is partially supported by PRIN 2017SSNZAW and PRIN 2022L34E7W, Moduli spaces and birational geometry. AT is partially supported by PRIN 2022HPSNCR: Semiabelian varieties, Galois representations and related Diophantine problems and PRIN 2020KKWT53: Curves, Ricci flat Varieties and their Interactions, and is a member of the INDAM group GNSAGA.
@@ 2.1. Resolving unibranch points
For brevity we write: “$q\in C$ is an $(m,n)$-point w.r.t. $F$”.
@@
Let $q\in C$ be an $(m,n)$-point w.r.t. $L$, where $L$ is a line locally at $q$; set $k=\lfloor n/m\rfloor$. Then $q^{h}\in C^{h}$ is an $(m,n-(h-1)m)$-point w.r.t. $L^{h}$ for every $h\leq k-1$, and $q^{k}$ is an $m$-fold point. Moreover
@@
    if $n/m\neq k$ then $q^{k}\in C^{k}$ is an $(m,n-(k-1)m)$-point w.r.t. $L^{k}$, and $q^{k+1}\in C^{k+1}$ is an $(n-km,m)$-point w.r.t. $E_{k+1}$;
@@
If $h\leq k-1$, the first inequality is strict, hence $n-(h-1)m>m$. Therefore $f_{h}(x_{h},y_{h})$ contains the summand $y_{h}^{m}$ but not the summand $x_{h}^{m}$. Hence, as $q^{h}\in C^{h}$ is unibranch, all terms of degree at most $m$ divisible by $x_{h}y_{h}$ must vanish. Hence ${\operatorname{mult}}_{q^{h}}(C^{h})=m$ and, as $y_{h}=0$ is the local equation of $L^{h}$, we get $(L^{h}\cdot C^{h})_{q^{h}}=n-(h-1)m$, hence $q^{h}\in C^{h}$ is an $(m,n-(h-1)m)$-point w.r.t. $L^{h}$, as claimed.
Let $h=k$. If $n/m\neq k$ then $k<n/m$ and the second inequality of (3) is strict. Hence, by the above argument, $q^{k}\in C^{k}$ is an $(m,n-(k-1)m)$-point w.r.t. $L^{k}$. Now look at $q^{k+1}\in C^{k+1}$, we have
@@
As before, $y_{k+1}^{m}$ is the smallest power of $y_{k+1}$ and $x_{k+1}^{n-km}$ the smallest power of $x_{k+1}$, but now $n-km<m$. Since $q^{k+1}$ is unibranch, by the same argument, ${\operatorname{mult}}_{q^{k+1}}(C^{k+1})=n-km$ and, as $x_{k+1}=0$ is the local equation of $E_{k+1}$, we conclude that $q^{k+1}\in C^{k+1}$ is an $(n-km,m)$-point w.r.t. $E_{k+1}$.
@@
If $k=1$ the first part of the statement is vacuous and the only interesting part is (a), which states that $q^{2}\in C^{2}$ is an $(n-m,m)$-point with respect to the exceptional divisor $E_{2}$
If $q\in C\subset\mathbb{P}^{2}$ is a singular unibranch point, hence an $(m,n)$-point, one can resolve the singularity by a finite sequence of blow-ups of $\mathbb{P}^{2}$, each centered at a point lying over $q$. This process leads to consider the euclidean algorithm for the pair $(m,n)$. This connection is known; see [Wal04], for example. Our setting here is different, as we need to extend our analysis to smooth points of curves and to tangent lines.
@@
where $l_{j}$ is defined inductively for $12\leq j\leq r+$ 1
@@
As it turns out, in the resolution sequence resolving the $(m,n)$-point $q$, the first multiplicities that appear over $q$ are $l_{1},l_{2},\ldots,l_{r}$, in this order. Moreover, each multiplicity $l_{j}$ appears exactly $k_{j}$ consecutive times.
We want to use this to give a convenient description of the resolution sequence. We define a total order on the following set of pairs associated to the euclidean sequence (4)
@@
Assume $n/m\not\in\mathbb{N}$, i.e. $k_{1}\neq n/m$. By Lemma 2.1.2, for every $i=1,\ldots,k_{1}$ the point $q^{i}_{1}\in C_{1}^{i}$ is an $(m,n-(i-1)m)$-point w.r.t. the proper transform, $F_{1}^{i}$, of $E^{1}_{1}$; moreover, $q^{1}_{2}\in C^{1}_{2}$ is an $(l_{2},l_{1})$-point with respect to $E^{1}_{2}$. Now we continue inductively on $j$; assume the statement for the level $j-1$ and consider the $j$-th level of the chain, with $12\leq j\leq r-$ 1. Then $q^{1}_{j}\in C^{1}_{j}$ is an $(l_{j},l_{j-1})$-point w.r.t. $E^{1}_{j}=F^{1}_{j}$, and $E^{1}_{j}$ is a line locally at $q^{1}_{j}$. By the same lemma, for every $i=1,\ldots,k_{j}$ the point $q^{i}_{j}\in C_{j}^{i}$ is an $(l_{j},l_{j-1}-(i-1)l_{j})$-point with respect to the proper transform, $F_{j}^{i}$, of $E^{1}_{j}$.
@@ 2.2. Normal crossings resolutions
Let $q\in C\subset\mathbb{P}^{2}$ be an $(m,n)$-point and consider the sequence of maps in Proposition 2.1.4. We add the blow-up of the last surface, $S_{r}^{k_{r}}$, at $q_{r}^{k_{r}}$, written
$\sigma_{*}:S_{*}\longrightarrow S_{r}^{k_{r}},$
let $C_{*}\subset S_{*}$ be the proper transform of $C$ and $q_{*}\in C_{*}$ the point lying over $q$. We set $\beta_{*}=\beta\circ\sigma_{*}$ so that we have, abusing notation,
@@
The subscript “” indicates that, if $c>1$, the type of singularity of $q_{*}$ is not known a priori, in which case $\beta_{*}$ is only a partial resolution of $q$. We write
(5) $D_{*}=D_{*}(q):=(\beta_{*}^{-1}(C\cup L))_{\operatorname{red}}$
for the reduced subscheme/divisor supporting $\beta_{*}^{-1}(C\cup L)\subset S_{*}$.
@@
    For every $(j,i)\not\in\{(r,k_{r}),(1,k_{r})\}$ there is a unique $(j^{\prime},i^{\prime})>(j,i)$ such that $A^{i}_{j}\cap A^{i^{\prime}}_{j^{\prime}}\neq\emptyset$, moreover
@@ 3.1. Dual graphs of partial resolutions
Let $C\subset\mathbb{P}^{2}$ have a unique (for simplicity) singular point in $q$, it is well known that there exists a unique surface $S_{\rho}$ with a birational morphism $\rho:S_{\rho}\to\mathbb{P}^{2}$ such that the strict transform of $C$ in $S_{\rho}$ is smooth, the reduced scheme underlying $\rho^{-1}(C)$ is a divisor with normal crossings, written $D_{\rho}$, and $S_{\rho}$ is minimal with respect to these properties. The dual graph, $G_{\rho}$, of $D_{\rho}$ (whose vertices are the irreducible components of $D_{\rho}$, with an edge between two vertices if the two components intersect) is often called the resolution graph of $q$; [Wal04]. The graph $G_{\rho}$ is not suitable here, and we will define a variant of it which will lead, in turn, to a tropical curve.
@@
of the divisor $D_{*}$, defined in Section 2.2. The vertices of $G_{*}$ are the irreducible components of $D_{*}$, hence $G_{*}$ has $2+\sum_{i=1}^{r}k_{i}$ vertices, and there is an edge between two vertices if the two components intersect. For the next lemma, recall that a leaf of a graph is a vertex of degree $1$ (and the degree, or valency, of a vertex is the number of half-edges adjacent to it).
@@
Let now $q^{\prime}$ be a $(2,8)$-point, then $q$ is singular, $r=1$ and $k_{1}=4$, and the graph $G(D_{*}(q^{\prime}))$ is the same as above.
@@
$C_{*}\longrightarrow C^{2}_{r}\longrightarrow C^{1}_{r}\longrightarrow C^{1}_{r-1}\longrightarrow\ldots\longrightarrow C^{1}_{2}\longrightarrow C$
and the graph $G_{*}$ is as follows, if $r$ is even.
@@
We ask under what conditions different pairs give the same graph.
@@ 3.2. The contact tropical curve of an $(m,n)$-point
$E(G_{q})=\left\{\begin{array}[]{ll}\{C_{*}E_{*},\ E_{*}A^{1}_{r},\ E_{*}A^{1}_{r-1},\ A^{1}_{j+1}A^{1}_{j-1},\ \forall j=2,\ldots,r-1\}&\ \text{if }r>1\\ \{C_{*}E_{*},\ E_{*}A^{1}_{1}\}&\ \text{if }r=1\\ \end{array}\right.$
@@
(9) $\ell_{q}(vw):=\left\{\begin{array}[]{ll}1&\ \text{if }vw=E_{*}C_{*}\text{ or }vw=E_{*}A^{1}_{r}\\ k_{r}&\ \text{if }vw=E_{*}A^{1}_{r-1}\\ k_{j}&\ \text{if }vw=A^{1}_{j+1}A^{1}_{j-1},\quad\forall j=2,\ldots,r-2.\\ \end{array}\right.$
@@
If $(m,n)\sim(m^{\prime},n^{\prime})$ then Lemma 3.1.5 yields $G^{(m,n)}_{*}\cong G^{(m^{\prime},n^{\prime})}_{*}$, hence $\Gamma^{(m,n)}\cong\Gamma^{(m^{\prime},n^{\prime})}$.
@@
Therefore $\Gamma$ and $\Gamma^{\prime}$ have the same number of vertices, equal to $r+2$, hence $r=r^{\prime}$. Moreover, $\Gamma$ and $\Gamma^{\prime}$ have the same number of legs, equal to $\min\{1,k_{1}-1\}=\min\{1,k^{\prime}_{1}-1\}$; if this number is 00 then $k_{1}=k^{\prime}_{1}=1$. If $\min\{1,k_{1}-1\}=1$ then $\Gamma$ and $\Gamma^{\prime}$ have each one leg of respective lengths $k_{1}-1$ and $k^{\prime}_{1}-1$. Since the length of the leg is preserved by the isomorphism $\phi$, we get $k_{1}=k^{\prime}_{1}$. We write

### v3

@@ 1.1. Results
In this paper we study unibranch points of curves, i.e. points which have a single preimage in the normalization of the curve. Our first goal is to study under what conditions two projective plane curves of degree at least $2$ intersect in exactly one point, and this point is unibranch for both of them. This problem is also relevant for the study of the Lang exceptional set appearing in some conjectures related to arithmetic and hyperbolic geometry, as we shall explain later in this introduction.
@@
Let $B,C\subset\mathbb{P}^{2}$ be integral curves of degree at least $2$ such that $B\cap C=\{q\}$; assume that $q$ is an $(m^{\prime},n^{\prime})$-point for $B$ and an $(m,n)$-point for $C$. Then there is an isomorphism of tropical curves between $\Gamma^{(m,n)}$ and $\Gamma^{(m^{\prime},n^{\prime})}$; equivalently, $m^{\prime}/n^{\prime}=m/n$.
@@ 1.2. Context
Our work in this paper is motivated by the algebro-geometric counterparts of some outstanding conjectures from diophantine and hyperbolic geometry. To introduce them, let $B\subset\mathbb{P}^{2}$ be a complex projective curve of degree at least $3$, having at most nodal singularities. The so-called algebraic exceptional set of $B$, introduced by Lang, is denoted as follows
@@
As an example of application of this type of results we mention that, in our previous paper [CT25], we used as a crucial ingredient a special case of Theorem 1.1.1 (Theorem 2.2.1 in loc.cit. which assumes $B$ is smooth at $q$) to show that $\mathcal{E}(B)$ is empty for a general $B$ with three irreducible components of total degree at least 5.
@@ 1.3. Notation
Given a point $p\in C$, we write ${\operatorname{mult}}_{p}(C)$ for the multiplicity of $C$ at $p$; if $m={\operatorname{mult}}_{p}(C)$ we say that $p$ is an $m$-fold point . We denote by $\nu_{C}:C^{\nu}\to C$ the normalization. If $|\nu_{C}^{-1}(p)|=1$ we say that $p$ is unibranch.
@@ Acknowledgements.
We are grateful to Pietro Corvaja, Kristin DeVleming, Luca Ferrigno, Nicola Ottolini, Wei Chen and an anonymous referee for useful comments. Special thanks to Laura Capuano for showing us Lemma 5.1.2. LC is partially supported by PRIN 2017SSNZAW and PRIN 2022L34E7W, Moduli spaces and birational geometry. AT is partially supported by PRIN 2022HPSNCR: Semiabelian varieties, Galois representations and related Diophantine problems and PRIN 2020KKWT53: Curves, Ricci flat Varieties and their Interactions, and is a member of the INDAM group GNSAGA.
@@ 2.1. Resolving unibranch points
For brevity we write: $q\in C$ is an $(m,n)$-point w.r.t. FF”.
@@
Let $q\in C$ be an $(m,n)$-point w.r.t. $L$, where $L$ is a line locally at $q$; set $k=\lfloor n/m\rfloor$. Then $q^{h}\in C^{h}$ is an (m,n(h1)m)(m,n-(h-1)m)-point w.r.t. $L^{h}$ for every $h\leq k-1$, and $q^{k}$ is an $m$-fold point. Moreover
@@
    if $n/m\neq k$ then $q^{k}\in C^{k}$ is an (m,n(k1)m)(m,n-(k-1)m)-point w.r.t. $L^{k}$, and $q^{k+1}\in C^{k+1}$ is an (nkm,m)(n-km,m)-point w.r.t. $E_{k+1}$;
@@
If $h\leq k-1$, the first inequality is strict, hence n(h1)m>mn-(h-1)m>m. Therefore $f_{h}(x_{h},y_{h})$ contains the summand $y_{h}^{m}$ but not the summand $x_{h}^{m}$. Hence, as $q^{h}\in C^{h}$ is unibranch, all terms of degree at most $m$ divisible by $x_{h}y_{h}$ must vanish. Hence ${\operatorname{mult}}_{q^{h}}(C^{h})=m$ and, as $y_{h}=0$ is the local equation of $L^{h}$, we get $(L^{h}\cdot C^{h})_{q^{h}}=n-(h-1)m$, hence $q^{h}\in C^{h}$ is an (m,n(h1)m)(m,n-(h-1)m)-point w.r.t. $L^{h}$, as claimed.
Let $h=k$. If $n/m\neq k$ then $k<n/m$ and the second inequality of (3) is strict. Hence, by the above argument, $q^{k}\in C^{k}$ is an (m,n(k1)m)(m,n-(k-1)m)-point w.r.t. $L^{k}$. Now look at $q^{k+1}\in C^{k+1}$, we have
@@
As before, $y_{k+1}^{m}$ is the smallest power of $y_{k+1}$ and $x_{k+1}^{n-km}$ the smallest power of $x_{k+1}$, but now nkm<mn-km<m. Since $q^{k+1}$ is unibranch, by the same argument, ${\operatorname{mult}}_{q^{k+1}}(C^{k+1})=n-km$ and, as $x_{k+1}=0$ is the local equation of $E_{k+1}$, we conclude that $q^{k+1}\in C^{k+1}$ is an (nkm,m)(n-km,m)-point w.r.t. $E_{k+1}$.
@@
If $k=1$ the first part of the statement is vacuous and the only interesting part is (a), which states that $q^{2}\in C^{2}$ is an (nm,m)(n-m,m)-point with respect to the exceptional divisor $E_{2}$
If $q\in C\subset\mathbb{P}^{2}$ is a singular unibranch point, hence an $(m,n)$-point, one can resolve the singularity by a finite sequence of blow-ups of $\mathbb{P}^{2}$, each centered at a point lying over $q$, such as the one described before Lemma 2.1.2. This process leads to consider the euclidean algorithm for the pair $(m,n)$. This connection is known; see [Wal04], for example. Our setting here is different, as we need to extend our analysis to smooth points of curves and to tangent lines.
@@
where $l_{j}$ is defined inductively for $2\leq j\leq r+1$
@@
As we shall see in Proposition 2.1.4, in the resolution sequence resolving the $(m,n)$-point $q$, the first multiplicities that appear over $q$ are $l_{1},l_{2},\ldots,l_{r}$, in this order. Moreover, each multiplicity $l_{j}$ appears exactly $k_{j}$ consecutive times.
We want to use this to give a more convenient notation for the resolution sequence. We define a total order on the following set of pairs associated to the euclidean sequence (4)
@@
Assume $n/m\not\in\mathbb{N}$, i.e. $k_{1}\neq n/m$. By Lemma 2.1.2, for every $i=1,\ldots,k_{1}$ the point $q^{i}_{1}\in C_{1}^{i}$ is an (m,n(i1)m)(m,n-(i-1)m)-point w.r.t. the proper transform, $F_{1}^{i}$, of $E^{1}_{1}$; moreover, $q^{1}_{2}\in C^{1}_{2}$ is an $(l_{2},l_{1})$-point with respect to $E^{1}_{2}$. Now we continue inductively on $j$; assume the statement for the level j1j-1 and consider the $j$-th level of the chain, with $2\leq j\leq r-1$. Then $q^{1}_{j}\in C^{1}_{j}$ is an $(l_{j},l_{j-1})$-point w.r.t. $E^{1}_{j}=F^{1}_{j}$, and $E^{1}_{j}$ is a line locally at $q^{1}_{j}$. By the same lemma, for every $i=1,\ldots,k_{j}$ the point $q^{i}_{j}\in C_{j}^{i}$ is an $(l_{j},l_{j-1}-(i-1)l_{j})$-point with respect to the proper transform, $F_{j}^{i}$, of $E^{1}_{j}$.
@@ 2.2. Normal crossings resolutions
Let $q\in C\subset\mathbb{P}^{2}$ be an $(m,n)$-point and consider the sequence of maps in Proposition 2.1.4. We now add the blow-up of the last surface, $S_{r}^{k_{r}}$, at $q_{r}^{k_{r}}$, written
$\sigma_{*}:S_{*}\longrightarrow S_{r}^{k_{r}}.$
Its exceptional divisor will be written $E_{*}$; let $C_{*}\subset S_{*}$ be the proper transform of $C$ and $q_{*}\in C_{*}$ the point lying over $q$. We set $\beta_{*}=\beta\circ\sigma_{*}$ so that we have, abusing notation,
@@
The subscript “” indicates that, if $c=\gcd(m,n)>1$, the type of singularity of $q_{*}$ is not known a priori (we know it is a $(c,h)$-point form some $h$) in which case, by part (d) of Proposition 2.1.4, $\beta_{*}$ is only a partial resolution of $q$.
Example 2.2.1.
Consider a curve $C\subset\mathbb{P}^{2}$ with a $(3,5)$ point, $q$, with respect to the line $L$. The euclidean sequence is $5>3>2>1$ hence $r=3$, and $k_{1}=k_{2}=1$, and $k_{3}=2$. The picture below represents the preimage of $C\cup L$ under four consecutive blow-ups. The first three blow-ups are those described in Proposition 2.1.4, the last, introduced above, is such that the resulting curve has normal crossings singularities. The first blow-up gives a $(2,3)$-point whose tangent is the proper transform of $L$. The second blow-up gives a $(1,2)$-point whose tangent is the exceptional divisor.
For better clarity, the curve $C$ and its proper transforms are curvilinear and red. The line $L$ and its proper transforms are horizontal and blue. The exceptional divisor of each blow-up is a vertical segment, and its proper transforms are segments (neither vertical nor horizontal).
$S_{*}$$E_{*}$$S_{3}^{2}$$E_{3}^{2}$$S_{3}^{1}$$E_{3}^{1}$$S_{2}^{1}$$E_{2}^{1}$$\mathbb{P}^{2}$LLCCqq
This example is a Fibonacci point, described in general in example 3.1.3.
We write
(5) $D_{*}=D_{*}(q):=(\beta_{*}^{-1}(C\cup L))_{\operatorname{red}}\subset S_{*}$
for the reduced subscheme/divisor supporting $\beta_{*}^{-1}(C\cup L)$. In the above picture, $D_{*}$ is the first from the left.
@@
    For every $(j,i)\not\in\{(r,k_{r}),(r,1)\}$ there is a unique $(j^{\prime},i^{\prime})>(j,i)$ such that $A^{i}_{j}\cap A^{i^{\prime}}_{j^{\prime}}\neq\emptyset$, moreover
@@ 3.1. Dual graphs of partial resolutions
Let $C\subset\mathbb{P}^{2}$ have a unique (for simplicity) singular point in $q$. It is well known that there exists a unique surface $S_{\rho}$ with a birational morphism $\rho:S_{\rho}\to\mathbb{P}^{2}$ such that the strict transform of $C$ in $S_{\rho}$ is smooth, the reduced scheme underlying $\rho^{-1}(C)$ is a divisor with normal crossings, written $D_{\rho}$, and $S_{\rho}$ is minimal with respect to these properties. The dual graph, $G_{\rho}$, of $D_{\rho}$ (whose vertices are the irreducible components of $D_{\rho}$, with an edge between two vertices if the two components intersect) is often called the resolution graph of $q$; [Wal04]. The graph $G_{\rho}$ is not suitable here, and we will define a variant of it which will lead, in turn, to a tropical curve.
@@
of the divisor $D_{*}$, defined in Section 2.2. The vertices of $G_{*}$ are the irreducible components of $D_{*}$, hence $G_{*}$ has $2+\sum_{i=1}^{r}k_{i}$ vertices, and there is an edge between two vertices if the two components intersect. We shall abuse notation and denote the vertices and the components they represent by the same symbol. For the next lemma, recall that a leaf of a graph is a vertex of degree $1$ (and the degree, or valency, of a vertex is the number of half-edges adjacent to it).
@@
Let now $q^{\prime}$ be a $(2,8)$-point, then $q^{\prime}$ is singular, $r=1$ and $k_{1}=4$, and the graph $G(D_{*}(q^{\prime}))$ is the same as above.
@@
$C_{*}\longrightarrow C^{2}_{r}\longrightarrow C^{1}_{r}\longrightarrow C^{1}_{r-1}\longrightarrow\ldots\longrightarrow C^{1}_{2}\longrightarrow C.$
The graph $G_{*}$ for Example 2.2.1 is the following
@@
If $r$ is even, $G_{*}$ is as follows
[…]
We now study under what conditions different pairs give the same graph.
@@ 3.2. The contact tropical curve of an $(m,n)$-point
[…]
@@
[…]
@@
If $(m,n)\sim(m^{\prime},n^{\prime})$ then Lemma 3.1.5 yields $G^{(m,n)}_{*}\cong G^{(m^{\prime},n^{\prime})}_{*}$, hence, by what we saw above, $\Gamma^{(m,n)}\cong\Gamma^{(m^{\prime},n^{\prime})}$.
@@
Therefore $\Gamma$ and $\Gamma^{\prime}$ have the same number of vertices, equal to $r+2$, hence $r=r^{\prime}$. Moreover, $\Gamma$ and $\Gamma^{\prime}$ have the same number of legs, equal to $\min\{1,k_{1}-1\}=\min\{1,k^{\prime}_{1}-1\}$; if this number is 0 then $k_{1}=k^{\prime}_{1}=1$. If $\min\{1,k_{1}-1\}=1$ then $\Gamma$ and $\Gamma^{\prime}$ have each one leg of respective lengths $k_{1}-1$ and $k^{\prime}_{1}-1$. Since the length of the leg is preserved by the isomorphism $\phi$, we get $k_{1}=k^{\prime}_{1}$. We write